# First-order CP phase transition in two-flavor QCD at $\theta=\pi$ under electromagnetic scale anomaly via a Nambu-Jona-Lasinio description

### v1

@@ Abstract
We discuss the thermal CP phase transition in QCD at $\theta=\pi$ under a weak magnetic field background, where the electromagnetic scale anomaly gets significant. To explicitize, we work on a two-flavor Polyakov-loop Nambu-Jona-Lasinio model at $\theta=\pi$ in the mean field approximation, including the electromagnetic-scale anomaly term. We find that the thermal CP phase transition becomes first order and the strength of the first order gets more prominent as the magnetic field increases. The associated potential barrier is thermally created by the electromagnetic scale anomaly and gives rise to criticality due to the induced potential of a non-perturbative form $\sim\frac{|eB|^{3}}{f_{\pi}}\frac{|P|}{P^{2}+m_{0}^{2}}$, where $eB$ denotes the magnetic field strength; $P$ the CP order parameter, and $m_{0}$ the isospin-symmetric current-quark mass. The CP-broken deconfinement(-like) domain, $T^{\rm(CP)}_{c}>T^{\rm(dec)}_{\rm pc}$, gets wider as $eB$ increases.
@@ I Introduction
In a view of a cosmic magnetic field background at the QCD epoch as aforementioned, the magnetic field strength might be redshifted to be as small as or smaller than the dynamical quark mass scale $m_{\rm dyn}\sim 300-400$ MeV: $eB\lesssim m_{\rm dyn}^{2}$. In that case, it has recently been clarified that the electromagnetic-scale anomaly effect is not negligible in a weak magnetic field and is shown to have important implications for the chiral phase transition properties regarding the quark mass dependence in QCD with $\theta=0$ Kawaguchi:2021nsa ; YuanyuanWang:2022nds ; YuanyuanWang:2024zwf . This scale anomaly effect has never been examined along with nonzero $\theta$.
In this paper, we discuss the effect of the electromagnetic scale anomaly by explicitly working on a two-flavor NJL model at $\theta=\pi$ with the Polyakov loop extension (PNJL model) in the mean field approximation (MFA). We observe that the thermomagnetic correction part in the electromagnetic scale anomaly term develops a potential barrier around the origin of the direction of the CP order parameter. It turns out that this induced potential term is of nonperturbative form, which cannot be described by the Ginzburg-Landau description.
The phase transitions along the chiral and Polyakov loop directions are smooth crossover. The deconfinement crossover goes not much different from the case without the electromagnetic scale anomaly. The CP-broken deconfinement(-like) domain is thus emergent: $T^{\rm(CP)}_{c}>T^{\rm(dec)}_{\rm pc}$, and it gets enlarged with a larger size of $eB$.
II PNJL model with electromagnetic scale anomaly
We begin by introducing the two-flavor PNJL model with nonzero $\theta$ under a constant external magnetic field. The Lagrangian of this model is as follows:
@@ II PNJL model with electromagnetic scale anomaly
$+g_{d}\left\{e^{i\theta}\operatorname{det}\left[\bar{\psi}\left(1+\gamma_{5}\right)\psi\right]+e^{-i\theta}\operatorname{det}\left[\bar{\psi}\left(1-\gamma_{5}\right)\psi\right]\right\}-\mathcal{U}(\Phi,\bar{\Phi};T),$
where $\psi=(u,d)^{T}$ denotes the two-flavor quark field and the isospin-symmetric current-quark mass $m_{0}$. The covariant derivative $D_{\mu}=\partial_{\mu}-iqA^{EM}_{\mu}-iA_{\mu}$ contains the couplings of quarks to the gluon field $A_{\mu}$, and the external electromagnetic field $A^{EM}_{\mu}$, with its electromagnetic charge matrix given by $q=e\cdot\operatorname{diag}\{2/3,-1/3\}$. For simplicity, the magnetic field $B$ is set along the $z$-axis and is embedded in the electromagnetic gauge field of the form $A^{EM}_{\mu}=(0,0,Bx,0)$. The gluon field $A_{\mu}$ includes only the temporal component $A_{\mu}=\delta_{\mu}^{0}A_{0}=-iA_{4}$, where $A_{4}$ denotes the field in Euclidean space. The four-fermion interaction with the coupling $g_{s}$ has the full chiral $U(2)_{L}\times U(2)_{R}$ invariance, and $\tau_{a}(a=0,\ldots,3)$ denotes the $U(2)$ generators including $\tau_{0}={\bf 1}_{2\times 2}$. The term with the coupling $g_{d}$ is called the Kobayashi-Maskawa-’t Hooft (KMT) determinant term Kobayashi:1970ji ; Kobayashi:1971qz ; tHooft:1976rip ; tHooft:1976snw and preserves the $SU(2)_{L}\times SU(2)_{R}\times U(1)_{V}$ invariance, but breaks the $U(1)$ axial symmetry.
[…]
$\mathcal{U}(\Phi,\bar{\Phi};T)=T^{4}\left[-\frac{b_{2}(T)}{2}\bar{\Phi}\Phi-\frac{b_{3}}{6}\left(\bar{\Phi}^{3}+\Phi^{3}\right)+\frac{b_{4}}{4}(\bar{\Phi}\Phi)^{2}\right],$ (2)
where
$b_{2}(T)=a_{0}+a_{1}\left(\frac{T_{0}}{T}\right)+a_{2}\left(\frac{T_{0}}{T}\right)^{2}+a_{3}\left(\frac{T_{0}}{T}\right)^{3}.$ (3)
The parameter settings in the Polyakov potential are fixed by the pure Yang-Mills lattice simulation with the following values Meisinger:2001cq : $a_{0}=6.75$, $a_{1}=-1.95$, $a_{2}=2.625$, $a_{3}=-7.44$, $b_{3}=0.75$ and $b_{4}=7.5$. $T_{0}=0.27\,\mathrm{GeV}$ denotes the critical temperature of the deconfinement phase transition in pure Yang-Mills theory.
We introduce the three-dimensional momentum cutoff $\Lambda$ to regularize the PNJL model in the MFA. We refer to empirical hadron observables in the isospin symmetric limit at $eB=T=0$ Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs : the pion mass $m_{\pi}=0.1402\,\mathrm{GeV}$, the pion decay constant $f_{\pi}=0.0926\,\mathrm{GeV}$, and quark condensate $\langle\bar{u}u\rangle=\langle\bar{d}d\rangle=(-0.2415\,\mathrm{GeV})^{3}$. Then the model parameters can be determined as $m=0.006\,\mathrm{GeV}$, $\Lambda=0.59\,\mathrm{GeV}$, and for $g_{s}=2(1-c)G_{0}$ with $g_{d}=2cG_{0}$ and $c=0.2$ Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs , we have $G_{0}\Lambda^{2}=2.435$.
@@
$+g_{d}\left\{\operatorname{det}\left[\bar{\psi^{\prime}}\left(1+\gamma_{5}\right)\psi^{\prime}\right]+\operatorname{det}\left[\bar{\psi^{\prime}}\left(1-\gamma_{5}\right)\psi^{\prime}\right]\right\}-\mathcal{U}(\Phi,\bar{\Phi};T),$
@@
$-2g_{d}\left(S_{u}^{\prime}S_{d}^{\prime}-P_{u}^{\prime}P_{d}^{\prime}\right)-\mathcal{U}(\Phi,\bar{\Phi};T),$
@@
$\Omega\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)$ $=\frac{g_{s}}{2}\left[\left(S_{u}^{\prime}+S_{d}^{\prime}\right)^{2}+\left(P_{u}^{\prime}+P_{d}^{\prime}\right)^{2}\right]+2g_{d}\left(S_{u}^{\prime}S_{d}^{\prime}-P_{u}^{\prime}P_{d}^{\prime}\right)$ (12)
$-\sum_{f=u,d}\sum_{n=0}^{\infty}\alpha_{n}\int\frac{dp_{z}}{2\pi}\frac{\left|q_{f}B\right|}{2\pi}f_{\Lambda}\left(p_{z},n\right)\left[3E_{f,n}\right.$
$+T\ln\left(1+3\Phi e^{-\frac{E_{f,n}}{T}}+3\bar{\Phi}e^{-2\frac{E_{f,n}}{T}}+e^{-3\frac{E_{f,n}}{T}}\right)$
$\left.+T\ln\left(1+3\bar{\Phi}e^{-\frac{E_{f,n}}{T}}+3\Phi e^{-2\frac{E_{f,n}}{T}}+e^{-3\frac{E_{f,n}}{T}}\right)\right]+\mathcal{U}(\Phi,\bar{\Phi};T),$
@@
and $f_{\varphi}$ stands for the vacuum expectation value of this $\varphi$ at $T=eB=0$. What we are concerned about is in a weak field limit, $eB\ll m_{\rm dyn}^{2}$, as noted in the Introduction #1#1#1 When $eB$ exceeds the threshold strength $\gtrsim m_{\rm dyn}^{2}$, the electromagnetic scale anomaly gets started to suppress. This is due to the fact that above the threshold, the dynamics of quarks will be dominated by the lowest Landau level states, which are longitudinally polarized along the magnetic field direction. Since the electromagnetic scale anomaly develops along the transverse direction, its effect becomes mostly remarkable in the weak magnetic regime Kawaguchi:2021nsa .. Then the trace anomaly $T_{\mu}^{\mu}$ in Eq.(14) takes the form YuanyuanWang:2024zwf
@@
$\Omega_{tot}\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)=\Omega\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)+V_{\text{eff}}^{(\mathrm{Tad})}\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},T,eB\right).$ (19)
When $T$ and $eB$ are given with $\theta=\pi$, the order parameters of the chiral, CP, and deconfinement(-like) phase transitions can be determined by the following stationary conditions:
$\frac{\partial\Omega_{tot}\left(\theta=\pi,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)}{\partial S_{f}^{\prime}}=0,\quad\frac{\partial\Omega_{tot}\left(\theta=\pi,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)}{\partial P_{f}^{\prime}}=0,$ (20)
$\frac{\partial\Omega_{tot}\left(\theta=\pi,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)}{\partial\Phi}=0,\quad\frac{\partial\Omega_{tot}\left(\theta=\pi,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)}{\partial\bar{\Phi}}=0\,.$
@@ III CP phase transition at $\theta=\pi$
In the case with $eB\neq 0$, but without the tadpole contributions (see the left panel in Fig. 1), the order parameter for both the chiral and CP symmetries, $P$, undergoes the normal second-order phase transition at the criticality (around $\sim 300$ MeV). This phase transition is fairly insensitive to the presence of the weak $eB$ (See comparison with the curve for $eB=0$#3#3#3 The critical point of the CP phase transition at $eB=0$ observed in the current analysis is slightly different from what has been seen in the literature Sakai:2011gs ; Huang:2024nbd based on the same PNJL model with the same set of the input parameters. This is because of the different regularization scheme: the smooth cutoff regulator is used as done in Eq.(5) or not., although the overall size of $P$ gets slightly amplified by $eB\neq 0$ due to the magnetic catalysis for the chiral symmetry breaking, as has been observed in the literature Chatterjee:2014csa ; Bandyopadhyay:2019pml .
@@
Figure 2: The deformation of the thermomagnetic potential projected onto the $\alpha\propto P$ direction, for the case with the tadpole potential term, in comparison with the case without the tadpole term, at $\theta=\pi$, $eB=0.02\,{\rm GeV}^{2}$, for various values of $T$ around the criticality ($T=T_{c}\simeq 305.7$ MeV). For simplicity, we have taken the isospin symmetric limit, where $P^{\prime}_{u}=P^{\prime}_{d}=0$ (i.e. $\beta_{u}=\beta_{d}=\beta=m_{0}$), $S^{\prime}_{u}=S^{\prime}_{d}=S^{\prime}=-\alpha/(2(g_{s}+g_{d}))$, $q_{u}=q_{d}=(2/3)e$, and $\Phi=\bar{\Phi}\simeq 0.7$ around $T=300$ MeV. Potential values have been adjusted by dividing by a factor of $(0.1\,{\rm GeV})^{4}$, just for convenience.
To clarify the barrier generation mechanism, in the left panel of Fig. 3 we plot the tadpole term contributions at $\theta=\pi$ as a function of $\alpha_{u}$ in Eq.(11), for $T=200$ MeV and 300 MeV with $eB=0.02\,{\rm GeV}^{2}$ fixed. The potential has been normalized as $V_{\rm eff}^{\rm Tad}(0)=0$. Again, for simplicity, we have taken the isospin symmetric limit as above. A spike structure is observed indeed in the thermomagnetic part (blue curves), which is not substantially affected even in the total tadpole term (black curves) including the vacuum term at $T=0$ (red curve). This sort of a peak structure is not seen when $\theta=0$ along the $\alpha$ direction (right panel of Fig. 3), where the isospin symmetric limit reads $P_{u}=P_{d}=0$, i.e., $\beta_{u}=\beta_{d}=\beta=0$, and $S_{u}=S_{d}=S=-(\alpha-m_{0})/(2(g_{s}+g_{d}))$.
Figure 3: (Left panel): the breakdown of the tadpole-potential term-contributions in Eq.(14), as a function of the CP order parameter $\alpha(\theta=\pi)=-2P(g_{s}+g_{d})$ normalized by the vacuum value at $T=eB=0$. The plots are split into the full contribution (black curve), the vacuum part at $T=0$ (red curve), and the thermomagnetic part (blue curve), for $T=200$ MeV and 300 MeV, with $eB=0.02\,{\rm GeV}^{2}$ fixed. The isospin symmetric limit has been applied as described in the main text. (Right panel): the same for $\alpha(\theta=0)=-2S(g_{s}+g_{d})$. The potential values in both panels have been adjusted by multiplying by some factor as in Fig. 2.
This discrepancy can be understood in an analytic way. First of all, consider only the thermomagnetic term in Eq.(14) (i.e., the second term in Eq.(16)) at $\theta=\pi$, and take the isospin symmetric limit as above. Then the corresponding tadpole potential term goes like
@@
where $\alpha$ at $\theta=\pi$ is related to $P$ in the original base as $\alpha=-2(g_{s}+g_{d})P$, and
@@
This has a peak at $\alpha=m_{0}$, hence a potential barrier is built there, and its magnitude gets greater as $eB$ increases. The peak structure appears because the method of reducing the quark-thermal loop integral from three dimensions to just one dimension (along the $p_{z}$ direction) works well. In this process, the spatial momentum integral is replaced by the area related to the magnetic field ($eB$) and a sum over Landau levels. Near the critical point, the lowest Landau level dominates the contribution to the tadpole term. This allows the dependence on $\alpha$ to be determined mainly by the scaling of the one-dimensional loop momentum $p_{z}$, as shown in Eq.(21). If the scaling of the other spatial momenta ($p_{x},p_{y}$) or alternatively higher Landau levels were considered, it would introduce additional $\alpha$-dependence, potentially disrupting the barrier generation structure.
@@
The phase transitions along the chiral $P^{\prime}(=S)$ and Polyakov loop ($\Phi$) directions are smooth crossover, where $P^{\prime}(=S)\sim 0$ at any $T$. Those crossover transitions proceed not much different from the case without the electromagnetic scale anomaly, nor $eB$ Sakai:2011gs ; Huang:2024nbd at $\theta=\pi$. See Fig. 4 showing the pseudocriticality of the deconfinement $(\Phi=1)$ confinement $(\Phi=0)$ transition, in which we observe $T_{\rm pc}^{({\rm dec})}\simeq 258$ MeV. Thus, the presence of the electromagnetic scale anomaly supports the generation of the CP-broken deconfinement(-like) domain $T^{\rm(CP)}_{c}>T^{\rm(dec)}_{\rm pc}$. This domain gets wider as $eB$ increases up to $\sim m_{\rm dyn}^{2}$.
Figure 4: The plot of the Polyakov loop variable $\Phi$ as a function of $T$, for the same parameter choice as in Fig. 1. All the curves look like overlapped including the case without the electromagnetic scale anomaly or even $eB$, which implies that the deconfinement phase transition is overall fairly insensitive to $eB$ in the present model study.

### v3

@@ Abstract
We discuss the thermal CP phase transition in QCD at $\theta=\pi$ under a weak magnetic field background, where the electromagnetic scale anomaly gets significant. To explicitize, we work on a two-flavor Nambu-Jona-Lasinio model at $\theta=\pi$ in the mean field approximation, including the electromagnetic-scale anomaly term. We find that the thermal CP phase transition becomes first order and the strength of the first order gets more prominent as the magnetic field increases. The associated potential barrier is thermally created by the electromagnetic scale anomaly and gives rise to criticality due to the induced potential of a non-perturbative form $\sim\frac{|eB|^{3}}{f_{\pi}}\frac{|P|}{P^{2}+m_{0}^{2}}$, where $eB$ denotes the magnetic field strength; $P$ the CP order parameter, and $m_{0}$ the isospin-symmetric current-quark mass.
@@ I Introduction
In a view of a cosmic magnetic field background at the QCD epoch as aforementioned, the magnetic field strength might be redshifted to be as small as or smaller than the intrinsic scale of QCD, $\Lambda_{\rm QCD}\sim 200-400$ MeV: $eB\lesssim\Lambda_{\rm QCD}^{2}$. In that case, it has recently been clarified that the electromagnetic-scale anomaly effect is not negligible in a weak magnetic field and is shown to have important implications for the chiral phase transition properties regarding the quark mass dependence in QCD with $\theta=0$ Kawaguchi:2021nsa ; YuanyuanWang:2022nds ; YuanyuanWang:2024zwf . This scale anomaly effect has never been examined along with nonzero $\theta$.
In this paper, we discuss the effect of the electromagnetic scale anomaly by explicitly working on a two-flavor NJL model at $\theta=\pi$ in the mean field approximation (MFA). We observe that the thermomagnetic correction part in the electromagnetic scale anomaly term develops a potential barrier around the origin of the direction of the CP order parameter. The associated potential barrier is thermally created by the electromagnetic scale anomaly and gives rise to criticality due to the induced potential of a nonperturbative form, which cannot be described by the Ginzburg-Landau description. $\sim\frac{|eB|^{3}}{f_{\pi}}\frac{|P|}{P^{2}+m_{0}^{2}}$, where $eB$ denotes the magnetic field strength; $P$ the CP order parameter, and $m_{0}$ the isospin-symmetric current-quark mass.
II NJL model with electromagnetic scale anomaly
We begin by introducing the two-flavor NJL model with nonzero $\theta$ under a constant external magnetic field. The Lagrangian of this model is as follows:
@@ II NJL model with electromagnetic scale anomaly
$+g_{d}\left\{e^{i\theta}\operatorname{det}\left[\bar{\psi}\left(1+\gamma_{5}\right)\psi\right]+e^{-i\theta}\operatorname{det}\left[\bar{\psi}\left(1-\gamma_{5}\right)\psi\right]\right\}$
where $\psi=(u,d)^{T}$ denotes the two-flavor quark field and the isospin-symmetric current-quark mass $m_{0}$. The covariant derivative $D_{\mu}=\partial_{\mu}-iqA^{EM}_{\mu}$ contains the couplings of quarks to the external electromagnetic field $A^{EM}_{\mu}$, with the electromagnetic charge matrix given by $q=e\cdot\operatorname{diag}\{2/3,-1/3\}$. For simplicity, the magnetic field $B$ is set along the $z$-axis and is embedded in the electromagnetic gauge field of the form $A^{EM}_{\mu}=(0,0,Bx,0)$. The four-fermion interaction with the coupling $g_{s}$ keeps the full chiral $U(2)_{L}\times U(2)_{R}$ invariance, and $\tau_{a}(a=0,\ldots,3)$ denotes the $U(2)$ generators including $\tau_{0}={\bf 1}_{2\times 2}$. The term with the coupling $g_{d}$ is called the Kobayashi-Maskawa-’t Hooft (KMT) determinant term Kobayashi:1970ji ; Kobayashi:1971qz ; tHooft:1976rip ; tHooft:1976snw and preserves the $SU(2)_{L}\times SU(2)_{R}\times U(1)_{V}$ invariance, but breaks the $U(1)$ axial symmetry.
We introduce the three-dimensional momentum cutoff $\Lambda$ to regularize the NJL model in the MFA. We refer to empirical hadron observables in the isospin symmetric limit at $eB=T=0$ Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs : the pion mass $m_{\pi}=0.1402\,\mathrm{GeV}$, the pion decay constant $f_{\pi}=0.0926\,\mathrm{GeV}$, and quark condensate $\langle\bar{u}u\rangle=\langle\bar{d}d\rangle=(-0.2415\,\mathrm{GeV})^{3}$. Then the model parameters can be determined as $m=0.006\,\mathrm{GeV}$, $\Lambda=0.59\,\mathrm{GeV}$, and for $g_{s}=2(1-c)G_{0}$ with $g_{d}=2cG_{0}$ and $c=0.2$ Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs , we have $G_{0}\Lambda^{2}=2.435$.
@@
$+g_{d}\left\{\operatorname{det}\left[\bar{\psi^{\prime}}\left(1+\gamma_{5}\right)\psi^{\prime}\right]+\operatorname{det}\left[\bar{\psi^{\prime}}\left(1-\gamma_{5}\right)\psi^{\prime}\right]\right\}$
@@
$-2g_{d}\left(S_{u}^{\prime}S_{d}^{\prime}-P_{u}^{\prime}P_{d}^{\prime}\right)$
@@
$\Omega\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},T,eB\right)$ $=\frac{g_{s}}{2}\left[\left(S_{u}^{\prime}+S_{d}^{\prime}\right)^{2}+\left(P_{u}^{\prime}+P_{d}^{\prime}\right)^{2}\right]+2g_{d}\left(S_{u}^{\prime}S_{d}^{\prime}-P_{u}^{\prime}P_{d}^{\prime}\right)$ (10)
$-N_{c}\sum_{f=u,d}\sum_{n=0}^{\infty}\alpha_{n}\int\frac{dp_{z}}{2\pi}\frac{\left|q_{f}B\right|}{2\pi}f_{\Lambda}\left(p_{z},n\right)\left[E_{f,n}+2T\ln\left(1+e^{-\frac{E_{f,n}}{T}}\right)\right]$
@@
and $f_{\varphi}$ stands for the vacuum expectation value of this $\varphi$ at $T=eB=0$. What we are concerned about is in a weak field limit, $eB\ll{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Lambda_{\rm QCD}^{2}}$, as noted in the Introduction #1#1#1 When $eB$ exceeds the threshold strength $\gtrsim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Lambda_{\rm QCD}^{2}}$, the electromagnetic scale anomaly gets started to suppress. This is due to the fact that above the threshold, the dynamics of quarks will be dominated by the lowest Landau level states, which are longitudinally polarized along the magnetic field direction. Since the electromagnetic scale anomaly develops along the transverse direction, its effect becomes mostly remarkable in the weak magnetic regime Kawaguchi:2021nsa .. Then the trace anomaly $T_{\mu}^{\mu}$ in Eq.(12) takes the form YuanyuanWang:2024zwf
@@
$\Omega_{tot}\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},T,eB\right)=\Omega\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},T,eB\right)+V_{\text{eff }}^{(\mathrm{Tad})}\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},T,eB\right).$ (17)
When $T$ and $eB$ are given with $\theta=\pi$, the order parameters of the chiral and CP phase transitions can be determined by the following stationary conditions:
$\frac{\partial\Omega_{tot}\left(\theta=\pi,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},T,eB\right)}{\partial S_{f}^{\prime}}=0,\quad\frac{\partial\Omega_{tot}\left(\theta=\pi,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},T,eB\right)}{\partial P_{f}^{\prime}}=0,$ (18)
@@ III CP phase transition at $\theta=\pi$
In the case with $eB\neq 0$, but without the tadpole contributions (see the left panel in Fig. 1), the order parameter for both the chiral and CP symmetries, $P$, undergoes the normal second-order phase transition at the criticality (around $\sim 300$ MeV). This phase transition is fairly insensitive to the presence of the weak $eB$ (See comparison with the curve for $eB=0$#3#3#3 The critical point of the CP phase transition at $eB=0$ observed in the current analysis is slightly different from what has been seen in the literature Sakai:2011gs ; Huang:2024nbd based on a similar NJL model description with the same set of the input parameters. This is because of the different regularization scheme: the smooth cutoff regulator is used as done in Eq.(3) or not., although the overall size of $P$ gets slightly amplified by $eB\neq 0$ due to the magnetic catalysis for the chiral symmetry breaking, as has been observed in the literature Chatterjee:2014csa ; Bandyopadhyay:2019pml .
@@
Figure 2: The deformation of the thermomagnetic potential projected onto the $\alpha\propto P$ direction, for the case with the tadpole potential term (black curves), in comparison with the case without the tadpole term (red-dashed curve at $T=320$ MeV), at $\theta=\pi$, $eB=0.02\,{\rm GeV}^{2}$, for various values of $T$ around the criticality ($T=T_{c}\simeq 320$ MeV). The dashed, solid, dotted black curves correspond to the cases with $T=300$ MeV, $320$ MeV, and $330$ MeV, respectively. For simplicity, we have taken the isospin symmetric limit, where $P^{\prime}_{u}=P^{\prime}_{d}=0$ (i.e. $\beta_{u}=\beta_{d}=\beta=m_{0}$), $S^{\prime}_{u}=S^{\prime}_{d}=S^{\prime}=-\alpha/(2(g_{s}+g_{d}))$, $q_{u}=q_{d}=(2/3)e$. Potential values have been adjusted by dividing by a factor of $(0.1\,{\rm GeV})^{4}$.
To clarify the barrier generation mechanism, in the left panel of Fig. 3 we plot the tadpole term contributions at $\theta=\pi$ as a function of $\alpha_{u}$ in Eq.(9), for $T=200$ MeV and 300 MeV with $eB=0.02\,{\rm GeV}^{2}$ fixed. The potential has been normalized as $V_{\rm eff}^{\rm Tad}(0)=0$. Again, for simplicity, we have taken the isospin symmetric limit as above. A spike structure is observed indeed in the thermomagnetic part (blue curves), which is not substantially affected even in the total tadpole term (black curves) including the vacuum term at $T=0$ (red curve). This sort of a peak structure is not seen when $\theta=0$ along the $\alpha$ direction (right panel of Fig. 3), where the isospin-symmetric limit gives $P_{u}=P_{d}=0$, i.e., $\beta_{u}=\beta_{d}=\beta=0$, and $S_{u}=S_{d}=S=-(\alpha-m_{0})/(2(g_{s}+g_{d}))$.
Figure 3: (Left panel): the breakdown of the tadpole-potential term-contributions in Eq.(12), as a function of the CP order parameter $\alpha(\theta=\pi)=-2P(g_{s}+g_{d})$ normalized by the vacuum value at $T=eB=0$. The plots are split into the full contribution (black curve), the vacuum part at $T=0$ (red curve), and the thermomagnetic part (blue curve), for $T=200$ MeV and 320 MeV, with $eB=0.02\,{\rm GeV}^{2}$ fixed. The isospin symmetric limit has been applied as described in the main text. (Right panel): the same for $\alpha(\theta=0)=-2S(g_{s}+g_{d})$. The potential values in both panels have been adjusted by multiplying by some factor as in Fig. 2.
This discrepancy can be understood in an analytic way. First of all, consider only the thermomagnetic term in Eq.(12) (i.e., the second term in Eq.(14)) at $\theta=\pi$, and take the isospin symmetric limit as above. Then the corresponding tadpole potential term goes like #4#4#4 With things written in terms of $\alpha$ as in Eq.(19), the tadpole potential term is free from the model parameter dependence, except $m_{0}$.
@@
up to a factor of $(N_{f}\cdot N_{c})=2\cdot 3$, where $\alpha$ at $\theta=\pi$ is related to $P$ in the original base as $\alpha=-2(g_{s}+g_{d})P$, and
@@
This has a peak at $\alpha=m_{0}$, hence a potential barrier is built there, and its magnitude gets greater as $eB$ increases. Note also that Eq.(22) is ultraviolet free, hence the emergence of the barrier, i.e., the induced first-order feature is free from the regularization scheme #5#5#5 In the literature Avancini:2019wed , it has been discussed that in the case of the NJL model similar to the present one, at $T=0$ an $eB$-dependent regulator as in Eq.(3) leads to an artificial wavy or oscillating behavior of the quark condensate in a strong $eB$ regime ($eB\gtrsim 0.2\,{\rm GeV}^{2}$). We have checked that for the weak $eB$ regime that we are currently concerned about, the CP order parameter $P$ at $T=0$ does not exhibit such a wavy trend. .
The peak structure is generated when the method of reducing the quark-thermal loop integral from three dimensions to just one dimension (along the $p_{z}$ direction) works well. Recall that as in Eq.(2), the spatial momentum integral in Eq.(19) has been replaced by the area related to the magnetic field ($eB$) and a sum over Landau levels. Since near the critical point, the lowest Landau level is dominant in the tadpole term, the $\alpha$ dependence arises in Eq.(19) mainly by the scaling of the one-dimensional loop momentum $p_{z}$. If the scaling of the other spatial momenta ($p_{x},p_{y}$) or alternatively higher Landau levels were considered, it would introduce additional $\alpha$-dependence, potentially disrupting the barrier generation structure.